# 4D Radar-Inertial Odometry based on Gaussian Modeling and Multi-Hypothesis Scan Matching

### v3

@@ VI-B Results on Snail-Radar
Table III shows results using the ARS548 radar. This radar produces much sparser point clouds than Oculii Eagle (around 10x fewer points). We believe this to be the effect of Oculii Eagle’s scan enhancement algorithm. However, we can observe better metrics in general compared to Oculii Eagle. The sparser scans trigger a known issue in PCL’s implementation of NDT that cause a program termination, so unfortunately NDT is excluded from the comparison. The results once again show different methods leading in different sequences, however it is worth pointing out that 6 out of the 10 best results are now obtained by our Gaussian algorithms, as well as 3 out of the 10 second best results. Another thing to mention is that the multi-hypothesis scan matching seems to perform worse than the single-hypothesis version. We hypothesize this to be related to the increased sparsity of the scans, and believe it could be solved by improving the sampling distribution of the particles.
VI-C Summary of results

### v4

@@ VI-B Results on Snail-Radar
Table II also shows results using the ARS548 radar. This radar produces much sparser point clouds than Oculii Eagle (around 10x fewer points). We believe this to be the effect of Oculii Eagle’s scan enhancement algorithm. However, we can observe better metrics in general compared to Oculii Eagle. The sparser scans trigger a known issue in PCL’s implementation of NDT that cause a program termination, so unfortunately NDT is excluded from the comparison. The results once again show different methods leading in different sequences, however it is worth pointing out that 6 out of the 10 best results are now obtained by our Gaussian algorithms, as well as 3 out of the 10 second best results. Another thing to mention is that the multi-hypothesis scan matching seems to perform worse than the single-hypothesis version. We hypothesize this to be related to the increased sparsity of the scans, and believe it could be solved by improving the sampling distribution of the particles.
VI-C Additional qualitative experiments
In order to further justify our modeling approach based on globally optimized 3D Gaussians versus existing voxel-based approaches, we compare in Fig. 2 the output of our method against NDT. To do this, we prepare a radar-based map of NTU4DRadLM’s cp sequence using the supplied ground truth, and model a section of said map with both NDT and our method. The positions and sizes of the Gaussians generated by NDT are clearly constrained and influenced by the voxel grid, while our approach generates better fitted Gaussians. Moreover, our method is able to preserve geometric details such as corners, and allocates a higher budget of Gaussians to areas with higher point density, resulting in overall better representation of the geometry.
Additionally, we include a video supplement. In the video we show several Gaussian models of the same map created using different points-per-Gaussian ratios. It can be seen that, regardless of said hyperparameter, our modeling method is able to generate a representation reasonably fitting the map. Different number of Gaussians are allocated to areas of the map depending on the density and complexity of the area, as well as the available budget. Moreover, when the ratio is higher (meaning fewer Gaussians), sparse regions with few (outlier) radar points do not receive any Gaussians. Finally, we also include an example of the radar-inertial odometry pipeline on the same NTU4DRadLM cp sequence, compared to the ground truth trajectory.
Figure 2: Qualitative comparison between NDT and our Gaussian modeling using NTU4DRadLM’s cp sequence.
VI-D Summary of results